# Prospects for detecting asteroid-mass primordial black holes in extreme-mass-ratio inspirals with continuous gravitational waves

### v1

@@ VII Sensitivity of NEMO to asteroid-mass PBHs
High-frequency gravitational-wave (HFGW) detectors may deliver exquisite sensitivity in frequency ranges currently lacking by current and proposed ground-based detectors Aggarwal et al. (2021). In terms of inspiraling PBHs, at such frequencies, the system will be extremely close to merger, and thus the purely CW approach discussed in Section III would break down. However, searches for “transient CWs” have been shown to be able to track rapid frequency variations of such inspiraling systems over time Miller et al. (2021b); Andrés-Carcasona et al. (2023); Alestas et al. (2024); Miller (2024); Miller et al. (2024a, b), and even in the case of “mini-EMRI” systems Guo and Miller (2022). While such searches, and also matched filtering ones, have not yet evolved to handle completely eccentric waveforms, it is worth considering how well such EMRI systems could be constrained in future HFGW detectors, in order to motivate the further development of these techniques.
NEMO is a planned high-frequency GW detector that plans to deliver exquisite sensitivity in the 2-4 kHz regime to increase the detection prospects of neutron star mergers, and is comparable to the sensitivities of ET and Cosmic Explorer. For our purposes, however, such high-frequency sensitivity simply implies that we could detect PBH inspirals that are closer to merger than those we have considered in Section V.
Assuming a rough power spectral density value of $S_{n}(f)=10^{-48}$ strain/Hz and a frequency range of [1000,2500] Hz, we calculate the constraint on $\tilde{f}$ that would arise from a future analysis of NEMO data. We compute the projected constraints following the formalism presented in Section IV; however, because the GW signals become more transient-like and less CW-like as $m_{2}$ increases, we actually compute the distance reach using Eq. 32 in Miller et al. (2021b). Unfortunately, we only achieve $\tilde{f}\lesssim 1$ for $m_{2}\gtrsim 10^{-8}M_{\odot}$ – see Fig. 10(a). Thus, we also compute the projected sensitivity assuming that we can use matched filtering. In this case, we calculate compute the luminosity distance reached following Maggiore (2008):
[…]
@@
×$\times\left(\frac{1-\left(f_{\text{min}}/f_{\text{max}}\right)^{4/3}}{7.053\times 10^{-1}}\right)\left(\frac{10^{-48}\text{ Hz}^{-1}}{S_{n}(f)}\right)$ (24)
In Fig. 10, we show this projected constraint for different choices of $m_{1}$. In the $m_{1}=0.1M_{\odot}$ curve, we observe a “kink” around $m_{2}\simeq 10^{-10}M_{\odot}$, which indicates a sort-of transition between when the signal duration exceeds the observation time, and when the signal becomes more “transient-like”. In practical terms, summing the contributions to the rate densities from different frequencies, outlined in Section IV, does not produce as competitive constraints as simply calculating the rate density in the Euclidean way:
@@
where $d_{\text{max}}=\text{max}_{f}[d(f)]$ and represents the system with a given $\mathcal{M}$ sweeping from $f_{\text{min}}$ to $f_{\text{max}}$ with a duration given by Eq. 6: $\Delta t<T_{\text{obs}}$, i.e. a “transient-like” signal, not a CW.
Figure 10: NEMO sensitivity between 1000 and 2500 Hz. The curves extend to different values of $m_{2}$ because we have fixed the chirp mass range to be $[10^{-10},10^{-4}]M_{\odot}$, which, for different choices of $m_{1}$, results in different values for $m_{2}$. Left: CW semi-coherent sensitivity. Right: matched-filter sensitivity. In the orange, green and red curves, we observe two “kinks”: one at $\sim 5\times 10^{-12}M_{\odot}$, and another at $\sim 10^{-10}M_{\odot}$. The first kink occurs when the signal duration no longer exceeds the observation time, i.e. it becomes more “transient-like”. In practical terms, summing the contributions to the rate densities from different frequencies, outlined in Section IV, does not produce as competitive constraints as simply calculating the rate density in the Euclidean way (Eq. 26). The second kink separates two regimes of different slopes, and results because, for sufficiently light $m_{2}$, $d_{\text{max}}$ occurs at a frequency between $[f_{\text{min}},f_{\text{max}}]$. This happens because signals with these frequencies, in the span of $T_{\text{obs}}$, would actually spin out of the $[f_{\text{min}},f_{\text{max}}]$ band analyzed. Thus, we limit the frequency that we allow to signal to spin up to to $f_{\text{max}}$, which corresponds to them lasting shorter than $T_{\text{obs}}$. As $m_{2}$ increases, the frequency that maximizes $d(f)$ shifts to lower and lower frequencies. When that frequency falls below 1000 Hz, we observe the second kink, and the slope changes, since after the second kink, the maximum distance reach is always obtained by analyzing the full frequency range.
@@ VIII Conclusions
In this paper, we have shown that CW and tCW searches could constrain the existence of asteroid-mass PBHs in EMRI systems currently and with future GW detectors. In the mass regime in which our projected constraints overlap with those from microlensing, our results provide complementary ways of probing PBHs that could have formed in binary systems, instead of isolated ones. Additionally, in the so-called “asteroid-mass” regime, our results indicate that GW detectors would provide the first-ever stringent constraints on the fraction of DM that PBHs could compose. We note that we have parameterized our constraints in terms of $\tilde{f}$ in order to remain model-agnostic, and therefore in order to directly and fairly compare with the $f_{\text{PBH}}$ limits that arise from microlensing experiments or theoretical considerations, we would have to know exactly which assumptions are made on the PBH mass functions and formation mechanisms.
@@
Our results show that CW search techniques, exactly as they are, could provide stringent constraints on $\tilde{f}$ in the ET (3G) era of GW detectors. Furthermore, if tCW methods are used now, they could provide tight constraints on $\tilde{f}$ using O3 and subsequent observing runs of advanced LIGO, Virgo and KAGRA, highlighting both the short-term and long-term impact of CW searches for asteroid-mass PBHs.
Appendix A Additional equations
We provide here more information about Eq. 8, in particular the parameter $\Lambda$ that encapsulates particular analysis choices made in a real search for quasi-monochromatic GWs. It can be written as Astone et al. (2014)
$\Lambda$ $=4.02\left(\frac{p_{0}(1-p_{0})}{\theta_{\text{thr}}^{2}p_{1}^{2}}\right)^{1/4}\times\sqrt{CR_{\text{thr}}-\sqrt{2}\mathrm{erfc}^{-1}(2\Gamma)}$
$p_{0}$ $=e^{-\theta_{\text{thr}}}-e^{-2\theta_{\text{thr}}}+\frac{1}{3}e^{-3\theta_{\text{thr}}}$
$p_{1}$ $=e^{-\theta_{\text{thr}}}-2e^{-2\theta_{\text{thr}}}+e^{-3\theta_{\text{thr}}}$ (27)
where, typically, $\theta_{\text{thr}}$ is the threshold to determine which time-frequency pixels in the equalized power spectrogram are important, $\Gamma$ is the chosen confidence level, and $CR_{\text{thr}}$ is a threshold on a detection statistic called the critical ratio, which corresponds to the number of standard deviations a particular outlier is away from the mean. When $\theta_{\text{thr}}=2.5$, $\Gamma=0.95$ and $CR_{\text{thr}}$=5, $\Lambda=12.81$.
Appendix B Dependence of constraints on $\delta f$
@@
We know that $\left\langle VT\right\rangle$ depends on $d_{\text{max}}^{3}$ and $T$:
$\left\langle VT\right\rangle=\frac{4}{3}\pi d(f)^{3}T.$ (28)
and we will work in the small $\delta f$ limit, that is $\delta f\ll f$. We can write an abbreviated form of Eq. 5:

### v2

@@ VII Sensitivity of NEMO to asteroid-mass PBHs
High-frequency gravitational-wave (HFGW) detectors can deliver exquisite sensitivity in frequency ranges currently lacking by current and proposed ground-based detectors Aggarwal:2020olq. In terms of inspiraling PBHs, at such frequencies, the system will be extremely close to merger, and thus the purely CW approach discussed in Section III would break down. However, searches for “transient CWs” have been shown to be able to track rapid frequency variations of such inspiraling systems over time Miller:2020kmv; Andrés-Carcasona:2023df; Alestas:2024ubs; Miller:2024rca; Miller:2024fpo; Miller:2024jpo, and even in the case of “mini-EMRI” systems Guo:2022sdd. While such searches, and also matched filtering ones, have not yet evolved to handle completely eccentric waveforms, it is worth considering how well such EMRI systems could be constrained in future HFGW detectors, in order to motivate the further development of these techniques.
NEMO Ackley:2020atn is a planned HFGW detector that plans to deliver exquisite sensitivity in the 2-4 kHz regime to increase the detection prospects of neutron star mergers, and is comparable to the sensitivities of ET and Cosmic Explorer. For our purposes, however, such high-frequency sensitivity simply implies that we could detect PBH inspirals that are closer to merger than those we have considered in Section V.
Assuming a rough power spectral density value of $S_{n}(f)=10^{-48}$ strain/Hz and a frequency range of [1000,2500] Hz, we calculate the constraint on $\tilde{f}$ that would arise from a future analysis of NEMO data. We compute the projected constraints following the formalism presented in Section IV; however, because the GW signals become more transient-like and less CW-like as $m_{2}$ increases, we actually compute the distance reach using Eq. 32 in Miller:2020kmv. Unfortunately, we only achieve $\tilde{f}\lesssim 1$ for $m_{2}\gtrsim 10^{-8}M_{\odot}$ – see Fig. 7(a). Thus, we also compute the projected sensitivity assuming that we can use fully-coherent matched filtering. In this case, we calculate the luminosity distance reached following maggiore2008gravitational:
$d=\frac{2}{5}\sqrt{\frac{5}{6}}\frac{c}{\pi^{2/3}}\left(\frac{G\mathcal{M}}{c^{3}}\right)^{5/6}\left(\int_{f_{\text{min}}}^{f_{\text{max}}}df\frac{f^{-7/3}}{S_{n}(f)}\right)^{1/2}\rho$ (21)
@@
$\times\left(\frac{1-\left(f_{\text{min}}/f_{\text{max}}\right)^{4/3}}{7.053\times 10^{-1}}\right)\left(\frac{10^{-48}\text{ Hz}^{-1}}{S_{n}(f)}\right)$ (22)
This represents an idealized scenario with a detection threshold on the signal-to-noise ratio $\rho=8$, consistent with many matched filtering searches. In Fig. 7, we show this projected constraint for different choices of $m_{1}$. We can observe two kinks in each curve in this plot. The first kink occurs when the signal duration no longer exceeds the observation time, i.e. it becomes more “transient-like” [^4] Here, “transient-like” means that a signal lasts for a duration less than $T_{\text{obs}}$: in particular, it could last as short as $\mathcal{O}($hours).. In practical terms, summing the contributions to the rate densities from different frequencies, outlined in Section IV, does not produce as competitive constraints as simply calculating the rate density in the Euclidean way (Eq. 24). The second kink separates two regimes of different slopes, and results because, for sufficiently light $m_{2}$, $d_{\text{max}}$ occurs at a frequency between $f_{\text{min}},f_{\text{max}}]$. This happens because signals with these frequencies, in the span of $T_{\text{obs}}$, would actually spin out of the $f_{\text{min}},f_{\text{max}}]$ band analyzed. Thus, we limit the frequency to which we allow to signal to spin up to $f_{\text{max}}$, which corresponds to them lasting shorter than $T_{\text{obs}}$. As $m_{2}$ increases, the frequency that maximizes $d(f)$ shifts to lower and lower frequencies. When that frequency falls below 1000 Hz, we observe the second kink, and the slope changes, since after the second kink, the maximum distance reach is always obtained by analyzing the full frequency range.
In practical terms, summing the contributions to the rate densities from different frequencies, outlined in Section IV, does not produce as competitive constraints as simply calculating the rate density in the Euclidean way:
@@
where $d_{\text{max}}$ and represents the system with a given $\mathcal{M}$ sweeping from $f_{\text{min}}$ to $f_{\text{max}}$ with a duration given by Eq. 6: $\Delta t<T_{\text{obs}}$, i.e. a “transient-like” signal, not a CW. For a chosen chirp mass, we calculate Eq. 22 over many possible $f_{\text{min}},f_{\text{max}}]$ windows, and select the maximum of those to use in Eq. 24.
We note that the major difference between Eq. 24 and Eq. 15 is that the former applies in the case of signals whose durations are less then $T_{\text{obs}}$, while the latter is applied for signals whose durations are greater than $T_{\text{obs}}$. As explained in Section IV, Eq. 15 allows us to account for the fact that multiple sources could be emitting GWs in the same frequency bin and thus enhance the signal strength.
Figure 7: Projected constraints on $\tilde{f}$ using a NEMO power spectral density between 1000 and 2500 Hz with two methodologies: (a) semi-coherent CW and (b) fully coherent matched filtering for different $m_{1}$. Coherent searches provide significantly stronger constraints at the expense of increased computational cost: CW methods can only constrain $\tilde{f}<1$ for $m_{2}>8\times 10^{-9}M_{\odot}$, while coherent matched filtering can constrain $\tilde{f}<1$ for $m_{2}>2\times 10^{-13}M_{\odot}$. In (b),two “kinks” appear: the first ($\sim 2\times 10^{-11}M_{\odot}$ for three of the curves, at $\sim 10^{-10}$ for the blue curve) when the signal duration becomes shorter than $T_{\text{obs}}$, and the second ($\sim 10^{-10}M_{\odot}$ for three of the curves, $\sim 10^{-9}$ for the blue curve) when the frequency maximizing $d(f)$ shifts below 1000 Hz – see text for details.
@@ VIII Conclusions
In this paper, we have shown that CW and transient CW searches could constrain the existence of asteroid-mass PBHs in EMRI systems currently and with future GW detectors. In the mass regime in which our projected constraints overlap with those from microlensing, our results provide complementary ways of probing PBHs that could have formed in binary systems, instead of isolated ones. Additionally, in the so-called “asteroid-mass” regime, our results indicate that GW detectors would provide the first-ever stringent constraints on the fraction of DM that PBHs could compose. We note that we have parameterized our constraints in terms of $\tilde{f}$ in order to remain model-agnostic, and therefore in order to directly and fairly compare with the $f_{\text{PBH}}$ limits that arise from microlensing experiments or theoretical considerations, we would have to know exactly which assumptions are made on the PBH mass functions and formation mechanisms.
@@
Our results show that CW search techniques, exactly as they are, could provide stringent constraints on $\tilde{f}$ in the ET era of GW detectors. However, future work will entail understanding to what extent tCW methods can address the nonlinear signal frequency evolution in Eq. 3. In this case, the signal may even last much shorter than, depending on the component masses and initial orbital separation, and thus requires an algorithm capable of efficiently tracking curved, non-stationary trajectories in the time-frequency plane. Multiple methods exist to perform this kind of time-frequency tracking Miller:2018rbg; Sun:2018hmm; Oliver:2018dpt; Banagiri:2019obu; Keitel:2019zhb; Miller:2020kmv; Miller:2024jpo; Alestas:2024ubs; however, their computational cost, sensitivity and sky coverage must be evaluated in order to assess to what extent they could be sensitive to asteroid-mass PBHs now and in next-generation GW interferometers.
Our results thus demonstrate that CW search techniques, when adapted to capture the distinctive frequency evolution of asteroid-mass PBH inspirals, can provide a powerful probe using both current and future GW observatories. This is in line with the adaption of CW methods to search for a variety of DM signatures DAntonio:2018sff; Pierce:2018xmy; Palomba:2019vxe; Grote:2019uvn; Guo:2019ker; Armaleo:2020efr; Vermeulen:2021epa; Miller:2022wxu; Manita:2023mnc; LIGOScientific:2021odm; KAGRA:2022dwb; Gottel:2024cfj – see Miller:2025yyx for a recent review. While next-generation detectors such as ET and NEMO will open up vast new regions of parameter space, even modest modifications to present-day search strategies could already yield meaningful constraints on the fraction of DM that PBHs can compose. Taken together, these findings highlight the untapped potential of CW methods to test scenarios for PBHs that are otherwise difficult to access, and they motivate the development of more flexible search pipelines capable of capturing the full diversity of long-lived, tCW signals.
Appendix A Dependence of constraints on $\delta f$
@@
We know that $\left\langle VT\right\rangle$ depends on $d_{\text{max}}^{3}$ and $T$ through Eq. 11, and we will work in the small $\delta f$ limit, that is $\delta f\ll f$. We can write an abbreviated form of Eq. 5: